# Shear viscosity of quark-gluon plasma at finite temperature and chemical potential and QCD phase transitions

### v1

@@ II.2 Shear viscosity in kinetic theroy with the relaxation time approximation
×$\times f_{j}^{(0)}\left(E_{j},T,\mu_{q}\right)\cdot v_{\mathrm{rel}}\cdot\sigma_{ij\rightarrow cd}\left(s,T,\mu_{q}\right).$
@@
The relative velocity $v_{\mathrm{rel}}$ can be calculated in the c.m frame with
$v_{\text{rel }}=\frac{\sqrt{\left(E_{i}^{*}E_{j}^{*}-\left|\mathbf{p}_{i}^{*}\right|\left|\mathbf{p}_{j}^{*}\right|\right)^{2}-\left(M_{i}M_{j}\right)^{2}}}{E_{i}^{*}E_{j}^{*}}$ (24)
@@
where the factor $(1-f^{(0)})$ represents the Pauli blocking for quarks and/or antiquarks in the final state; $t$ is the Mandelstam variable; $t^{\pm}$ indicates the upper and lower limits of $t$
@@
±$\!\pm\!2\!\left.\sqrt{\!\frac{\left(s\!+\!M_{i}^{2}\!-\!M_{j}^{2}\right)^{2}}{4s}\!-\!M_{i}^{2}}\right.\!\sqrt{\frac{\left(s\!+\!M_{c}^{2}\!-\!M_{d}^{2}\right)^{2}}{4s}\!-\!M_{c}^{2}}.$
@@
$\frac{1}{4N_{c}^{2}}|{\mathcal{M}}|^{2}$ denotes the matrix element squared averaged over the color and spin of the incident particles, and summed over the final scattered particles. The calculation of $|{\mathcal{M}}|^{2}$ is related to the specific scattering channels, which can be performed using the standard methods in quantum field theory. One can refer to Refs. Rehberg et al. (1996a); Soloveva et al. (2021) for more details.
For the quark-quark scattering, only $u$ and $t$ channel are involved. The mesons exchanged in different scattering processes are listed in Table. 1. On the premise of isospin symmetry of $u,d$ quark, the process of $u+s\rightarrow u+s$ is equivalent to that of $d+s\rightarrow d+s$, and the process of $u+u\rightarrow u+u$ is equivalent to that of $d+d\rightarrow d+d$. For the quark-antiquark scattering, only $t$ and $s$ channel are involved. The mesons exchanged in different processes are listed in Table. 2. The antiquark-antiquark scattering can be calculated using the same steps as the quark-quark scattering.
@@ III Numerical results and discussion
In this section, we present the numerical results, including the Mott dissociation of mesons, the cross section, the flavor dependent relaxation time and the shear viscosity to entropy density ratio in the full QCD diagram. In particular, we will discuss the dependence of the relaxation processes and the transport coefficient on temperature and chemical potential as well as the QCD phase transition.
III.1 Mott dissociation of mesons
The dissociation of a meson involves its physical mass and constitute mass in the QGP medium. When the physical mass of a bound meson equals the constituent one, the meson becomes unstable and undergoes a Mott dissociation (Mott phase transition) Blaschke et al. (2017). The corresponding temperature is referred to as the Mott phase transition temperature, $T_{\text{Mott}}$. Above the $T_{\text{Mott}}$, mesons will dissociate into quark-antiquark pairs. From the onset of Mott phase transition the meson decay width becomes non-zero and increases with rising temperature.
@@ III.1 Mott dissociation of mesons
Figure 1 presents the physical masses, constitute ones and decay widths of pseudoscalar mesons (pion and kaon) as functions of temperature at $\mu_{B}=0$ and $\mu_{B}=600$ MeV. It shows that the constitute mass $M_{u}+M_{d}$ ($M_{u}+M_{s}$) of pion (kaon) decreases with increasing temperature due to the partial restoration of chiral symmetry. However, the physical mass of bound pion (kaon) derived from Eqs. (14) and (15) increases with rising temperature. Fig. 1 also indicates that the Mott phase transition at $\mu_{B}=600$ MeV is lower than that at $\mu_{B}=0$. This is because that not only the temperature but also the density effect can cause the restoration of chrial symmetry. It can be also observed that the pion and kaon undergo the Mott dissociations at nearly the same but slightly different temperature.
Figure 2: Mott phase transitions of pion and kaon and chiral phase transition of light $u,\,d$ quarks in the $T-\mu_{B}$ plane. The isentropic trajectories for $s/\rho_{B}=300,100,50,25,15,10,5,4,3$ are also plotted for the convenience of the subsequent discussion.
Figure 3: Mott phase transitions of pion and kaon and chiral phase transition of light $u,\,d$ quark in the $T-\rho_{B}$ plane.
@@
To indicate the relationship between the Mott dissociation of pseudoscalar mesons and the chiral phase transition of quarks, we draw in Figs. 2 and 3 these phase transition lines in the full QCD phase diagram. Fig. 2 demonstrates that for each given chemical potential the temperature of Mott phase transition is close to but slightly higher than that of chiral crossover transition of $u$ ($d$) quark. The difference between them is less than $10$ MeV in the whole crossover region, and continually decrease when approaching the CEP. Although these phase transition lines are close to each other in the $T-\mu_{B}$ plane near the critical region, there is a larger difference in the temperature-density plane, as demonstrated in Fig. 3. This attributes to that the baryon density in the critical region being sensitive to the chemical potential. Fig. 3 also indicates that the Mott dissociation occurs in the metastable phase for pion and both the metastable and unstable phases for kaon in the first-order phase transition range at lower temperature. The isentropic trajectories are complicated in the spinodal region as shown in Fig. 2, which implies that the thermodynamic and transport properties of QGP are complex in the range of the first-order phase transition when the spinodal structure is considered.
III.2 Shear viscosity
We present in Fig. 4 the cross sections of quark-quark and quark-antiquark scatterings at different temperature as functions of $\sqrt{s}$ for $\mu_{B}=0$ and $\mu_{B}=800$ MeV. For each scattering process, it is found that the cross section is relatively larger at the temperatures near the crossover phase transition and Mott phase transition. When the temperature is much higher than that of Mott phase transition, the cross sections decrease since the $\pi$ and $K$ mesons have larger decay widths, as shown in Fig. 4. Fig. 4 also shows that the thresholds of particles in each process change with temperature and chemical potential. Compared with $ud\rightarrow ud$ ($u\bar{d}\rightarrow u\bar{d}$), the threshold of $us\rightarrow us$ ( $u\bar{s}\rightarrow u\bar{s}$) at the same temperature and chemical potential is postponed to a larger $\sqrt{s}$. The reason is that $\sqrt{s}$ should at least equal to the larger of the sum of incident particle mass and the sum of outgoing particle mass, and $M_{u}+M_{s}$ is always larger than $M_{u}+M_{d}$ in the whole QCD phase diagram. Fig. 4 also indicates that the cross sections of quark-antiquark scattering are much larger than those of quark-quark scattering. This mainly attributes to the resonance of exchanged meson with the incident quarks in the $s$ channel in quark-antiquark scattering, which leads to a distinct peak-like structure existing in Fig. 4(c) and (d).
@@ III.2 Shear viscosity
Figure 4: Cross sections of quark-quark (including $ud\rightarrow ud$ and $us\rightarrow us$ ) and quark-antiquark (including $u\bar{d}\rightarrow u\bar{d}$ and $u\bar{s}\rightarrow u\bar{s}$) elastic scattering as functions of center-of-mass collision energy at different temperature and quark chemical potential.
@@
Figure  5 present the relaxation time of different particle species as functions of temperature at $\mu_{B}=0$ and $\mu_{B}=800$ MeV. For the $u,d$ isospin symmetric quark matter, the relaxation time $\tau_{u}=\tau_{d}$ and $\tau_{\bar{u}}=\tau_{\bar{d}}$ always hold. In particular, $\tau_{u}=\tau_{d}=\tau_{\bar{u}}=\tau_{\bar{d}}$ at vanishing chemical potential. In the case of $\mu_{B}=0$ , it can be seen that the relaxation time $\tau_{s}$ of strange quark is larger than that of $u$ ($d$) quark near the region of chiral and mott phase transitions. The chiral crossover transition temperature of $s$ quark is about $250$ MeV at $\mu_{B}=0$. We can see that $\tau_{s}$ gradually approaches $\tau_{u}$ at high temperature after the chiral restoration of $s$ quark, which is just the expected behavior under the SU(3) symmetry. At finite chemical potential, $\tau_{u}\neq\tau_{\bar{u}}$ and $\tau_{s}\neq\tau_{\bar{s}}$ due to the different distribution function of quark and antiquark. Numerically, the relaxation time of quarks are larger than those of antiquarks. Fig. 5 also indicates that the relaxation for each particle species has a concave shape, which is to certain degree connected with the shear viscosity of QGP medium.
@@
Figure 5: Relaxation time of quarks and antiquarks as functions of temperature for $\mu_{B}=0$ and $\mu_{B}=800$ MeV. The solid (hollow) triangles mark the temperatures of chiral (Mott) phase transitions for different particle species.
@@
In the following, we elaborate on the shear viscosity of QGP under different conditions. Fig. 6 presents the shear viscosity to entropy density ratio $\eta/s$ as functions of temperature at different chemical potentials. For $\mu_{B}$=0, the value of $\eta/s$ are very small above the temperature of meson Mott phase transition. The minimum of $\eta/s$ is about 0.1, close to the KSS limit, which indicates that the QGP medium is a nearly perfect” fluid in the extreme relativistic high-energy collision. The value of $\eta/s$ increases with the rising chemical potential. It can be observed from Fig. 6 that $\eta/s$ is slightly larger than 1 near the CEP, and the value will continually increase at lower temperature for larger chemical potentials.
@@
Figure 6: Shear visicosity to entropy ratio $\eta/s$ as functions of temperature at different $\mu_{B}$. The solid (hollow) triangles mark the temperatures of chiral (Mott) phase transitions.
@@
The minimum of $\eta/s$ at $\mu_{B}=0$ is closer to the Mott phase transition (meson dissociation) than the chiral phase transition of $u,d$ quark, and $\eta/s$ increases rapidly with the occurrence of hadronization. In Fig. 6 some values of $\eta/s$ below the chiral phase line are also preserved to display the variation of $\eta/s$ near the chiral crossover phase transition region. However we should keep in mind that the collisions between hadrons dominate the transport process in the hadronic phase at low temperature. It can be observed from Fig. 6 that the location of the minimum $\eta/s$ at finite chemical potential become more and more flat, and the relationship with phase transitions is not as distinct as that of $\mu_{B}=0$.
@@
Fig. 7 describes the behaviors of $\eta/s$ as functions of chemical potential at several temperatures lower than or equal to 250 MeV. For the temperature higher than that of the chiral phase transition of u(d) quark, e.g., $T=220$ or $250$ MeV, it can be seen that $\eta/s$ monotonically increases with rising chemical potential. The nonmonotonic behavior appears with the decrease of temperature, and the location of the minimum of $\eta/s$ gradually approaches to the chiral phase transition (solid triangle in Fig. 7), particularly in the critical region. As indicated in the formula of shear viscosity in Eq. (21), the contribution of density effect becomes more and more important with the rising chemical potential and the decreasing temperature.
Figure 7: $\eta/s$ as functions of baryon chemical potential at different temperatures. The solid triangles (dots) mark the chemical potentials of chiral phase transition in the crossover region (in the first-order region).
@@
Fig. 7 also shows the features of $\eta/s$ at $T=100,75,50$ MeV with inclusion of the first-order phase transition associated with the spinodal structure. The solid dots on each curve with the same color indicates the value of $\eta/s$ on the boundaries of the first-order phase transition. From higher chemical potentials to lower ones, the first solid dot encountered on each curve corresponds to the high-density side of the first-order phase transition, and the second solid dot corresponds to the low-density side. This figure indicates that the values of $\eta/s$ are quite different on the boundaries of the chiral broken and chiral restored phases at lower temperature.
Figure 8: $\eta/s$ as functions of baryon density at different temperatures. The solid triangles (dots) mark the baryon densities of chiral phase transition in the crossover region (in the first-order transition region).
@@
Fig. 8 presents the value of $\eta/s$ in the $T-\rho_{B}$ plane for different temperatures. The dashed curve is the boundary of the first-order phase transition. The behaviors of $\eta/s$ at $T=200$ MeV or higher temperatures are similar to those in the $T-\mu_{B}$ plane plotted in Fig. 7. However, compared to Fig. 7, in the $T-\rho_{B}$ plane the distance between the density of the minimum $\eta/s$ and that of the chiral phase transition appears to be enlarged when approaching to the critical region of the CEP from the high-temperature side. The reason is that the baryon density is sensitive to chemical potential in the critical region. Small variation of chemical potential will lead to a larger change of baryon number density. For the first-order phase transition, the difference of $\eta/s$ on the two phases is small near the CEP. However, it can be seen that at lower temperature, e.g., $T=50$ MeV, $\eta/s$ in the chiral restored QGP phase is much larger than the chiral breaking phase. This figure clearly indicates the shear viscosity to entropy density ratio at high baryon densities highly depends on the temperature. The value of $\eta/s$ increases with the decline of temperature, which is qualitatively consistent with the variation of shear viscosity of common liquids.
@@
Figure 9: The value of $\eta/s$ on the phase boundaries including the chiral corssover and first-order phase transitions.

### v2

@@ II.2 Shear viscosity and electric conductivity in kinetic theroy with the relaxation time approximation
$\times f_{j}^{(0)}\left(E_{j},T,\mu_{q}\right)\cdot v_{\mathrm{rel}}\cdot\sigma_{ij\rightarrow cd}\left(s,T,\mu_{q}\right).$
@@
The relative velocity $v_{\mathrm{rel}}$ can be calculated in the c.m frame with [97, 73]
$v_{\text{rel }}=\frac{\sqrt{\left(E_{i}^{*}E_{j}^{*}-\mathbf{p}_{i}^{*}\cdot\mathbf{p}_{j}^{*}\right)^{2}-\left(M_{i}M_{j}\right)^{2}}}{E_{i}^{*}E_{j}^{*}}$ (25)
@@
where the factor $(1-f^{(0)})$ represents the Pauli blocking for quarks and/or antiquarks in the final state; $t$ is the Mandelstam variable; $t_{\pm}$ indicates the upper and lower limits of $t$ [73]
@@
$\!\pm\!2\!\left.\sqrt{\!\frac{\left(s\!+\!M_{i}^{2}\!-\!M_{j}^{2}\right)^{2}}{4s}\!-\!M_{i}^{2}}\right.\!\sqrt{\frac{\left(s\!+\!M_{c}^{2}\!-\!M_{d}^{2}\right)^{2}}{4s}\!-\!M_{c}^{2}}.$
@@
$\frac{1}{4N_{c}^{2}}\sum_{sc}|{\mathcal{M}}|^{2}$ denotes the matrix element squared averaged over the color and spin of the incident particles, and summed over the final scattered particles. The calculation of $|{\mathcal{M}}|^{2}$ is related to the specific scattering channels, which can be performed using the standard methods in quantum field theory. One can refer to Refs. [80, 73] for more details.
For the quark-quark scattering, only $u$ and $t$ channel are involved. The mesons exchanged in different scattering processes are listed in Table. 1. Due to the isospin symmetry of $u,d$ quark, the process of $u+s\rightarrow u+s$ is equivalent to that of $d+s\rightarrow d+s$, and the process of $u+u\rightarrow u+u$ is equivalent to that of $d+d\rightarrow d+d$. For the quark-antiquark scattering, only $t$ and $s$ channel are involved. The mesons exchanged in different processes are listed in Table. 2. The antiquark-antiquark scattering can be calculated using the same steps as the quark-quark scattering.
@@ III Numerical results and discussion
In this section, we will present the numerical results, including the meson Mott dissociation, the cross section, the flavor dependent relaxation time and the shear viscosity and electric conductivity across the QCD phase diagram. In particular, we will discuss how the relaxation processes and transport coefficients depend on temperature, chemical potential (net baryon density), and the various phase transitions.
III.1 Mott dissociation of pseudoscalar mesons
The dissociation of mesons is closely related to the microscopic scattering of particles in QCD medium. It involves the physical mass and constitute mass of meson in the medium. When the mass of a bound meson equals the constituent one, the meson becomes unstable and undergoes a Mott dissociation (Mott phase transition) [98]. The corresponding temperature is referred to as the Mott phase transition temperature, $T_{\text{Mott}}$. Above the $T_{\text{Mott}}$, mesons will dissociate into quark-antiquark pairs. From the onset of Mott phase transition the meson decay width becomes non-zero and increases with rising temperature.
Figure 1 presents the physical masses of pseudoscalar mesons (pion and kaon), the constitute masses and the decay width as functions of temperature at $\mu_{B}=0$ (upper) and $\mu_{B}=600\,$MeV (lower). It shows that the constitute mass $M_{u}+M_{d}$ ($M_{u}+M_{s}$) of pion (kaon) decreases with increasing temperature due to the partial restoration of chiral symmetry. However, the physical mass of bound pion (kaon) derived from Eqs. (14) and (15) increases with the rising temperature. Fig. 1 also indicates that the Mott phase transition at $\mu_{B}=600$ MeV is lower than that at $\mu_{B}=0$. This is because that not only the temperature but also the density effect can cause the restoration of chrial symmetry. It can be also observed that the pion and kaon undergo the Mott dissociations at nearly the same but slightly different temperature.
@@ III.1 Mott dissociation of pseudoscalar mesons
To indicate the relationship between the Mott phase transition of pseudoscalar mesons and the quark chiral phase transition, we plot in Figs. 2 and 3 these phase transition lines in the full QCD phase diagram. Fig. 2 demonstrates that for each given chemical potential the pion and kaon Mott phase transition temperatures are close to but slightly higher than that of chiral crossover transition of $u$ ($d$) quark. The difference between them is less than $10$ MeV in the whole crossover region, and continually decreases when approaching the CEP. Although these phase transition lines are close to each other in the $T-\mu_{B}$ plane near the critical region, there is a relatively larger difference in the temperature-density plane, as demonstrated in Fig. 3. This is attributed to that the net baryon density in the critical region being sensitive to the chemical potential.
Figure 2: QCD phase diagram in the $T-\mu_{B}$ plane in the PNJL model, including the Mott phase transitions of pion and kaon and chiral crossover and first-order phase transitions with spinodal lines. The CEP locates $T_{C}=131\,$MeV, $\mu_{C}=873\,$MeV. The isentropic trajectories for $s/\rho_{B}=300,100,50,25,15,10,5,4,3$ are also plotted for the convenience of subsequent discussions.
Figure 3: QCD phase diagram in the $T-\rho_{B}$ plane in the PNJL model, including the Mott phase transitions of pion and kaon and chiral crossover and first-order phase transitions with spinodal lines. The CEP locates $T_{C}=131\,$MeV, $\rho_{C}=1.94\rho_{0}$). The isentropic trajectories for $s/\rho_{B}=5,4,3$ are also illustrated.
For a first-order phase transition, it is always accompanied by the presence of unstable and metastable phases in thermodynamics. As shown in Fig. 3, the spinodal line represents the boundary of the unstable state. Within the spinodal line, the compressibility is negative with the mechanical instability of system. The binodal line (first-order phase transition line) is the boundary between metastable and stable states. The region between the spinodal line and the binodal line is the metastable phase. For a given temperature, the baryon chemical potential $\mu_{B}$ exhibits a non-monotonic dependence on density across the first-order phase transition—increasing, then decreasing, and finally increasing again—as the system passes through the metastable and unstable regions. The spinodal line is defined by the locus of points where the derivative $\partial\mu_{B}/\partial\rho_{B}$ vanishes [99]. Fig. 3 also indicates that the Mott dissociation occurs in the metastable phase for pion and both the metastable and unstable phases for kaon in the first-order phase transition range at lower temperatures.
For smaller values of $s/\rho_{B}$, for example, $s/\rho_{B}=3,4,5$, the isentropic trajectories cross the first-order phase transition line as well as the metastable and unstable regions. The corresponding spinodal sturcture and isentropic trajectories in the $T-\mu_{B}$ plane are also plotted in Fig. 2, which predicates the underlying complexity of the transport properties of QGP in the evolution near the first-order phase transition when the spinodal region is involved.
III.2 Flavor dependent relaxation time
@@
Figure 4: Cross sections of quark-quark ( $ud\rightarrow ud$ and $us\rightarrow us$) and quark-antiquark ( $u\bar{d}\rightarrow u\bar{d}$ and $u\bar{s}\rightarrow u\bar{s}$) elastic scattering as functions of center-of-mass collision energy at different temperature and quark chemical potential.
The relaxation time of quark and antiquark plays a crucial role on the transport properties of QCD medium in kinetic theory with relaxation time approximation, which is related to the cross sections among different particles.
We present in Fig. 4 the cross sections of the $2\rightarrow 2$ elastic scattering for quark-quark and quark-antiquark scatterings as functions of center-of-mass energy $\sqrt{s}$ at $\mu_{B}=0$ and $\mu_{B}=800$ MeV. For each scattering process, it is found that the cross section is relatively larger at the temperatures near the crossover phase transition and Mott phase transition. When the temperature is much higher than that of Mott phase transition, the cross sections decrease since the $\pi$ and $K$ mesons have larger decay widths, as shown in Fig. 1. Fig. 4 also shows that the thresholds of particles in each process change with temperature and chemical potential. Compared with the $ud\rightarrow ud$ ($u\bar{d}\rightarrow u\bar{d}$) scattering, the threshold of $us\rightarrow us$ ( $u\bar{s}\rightarrow u\bar{s}$) process at the same temperature and chemical potential is postponed to a larger $\sqrt{s}$. The reason is that $\sqrt{s}$ should at least equal to the larger of the sum of incident particle mass and the sum of outgoing particle mass, and $M_{u}+M_{s}$ is always larger than $M_{u}+M_{d}$ in the whole QCD phase diagram. Fig. 4 also displays that some peak values in quark-antiquark scattering are obviously lager than those in quark-quark scattering. In the $u\bar{d}{\rightarrow}u\bar{d}$ and $u\bar{s}{\rightarrow}u\bar{s}$ scattering, this phenomenon appears at $\mu_{B}=0$ MeV, $T=180$, $210$ and 250250\,MeV and $\mu_{B}=800\,$MeV, $T=150$, $160$ and 180180\,MeV. This is mainly attributed to the resonance of the exchanged meson with the incident quarks in the $s$ channel in quark-antiquark scattering, which leads to a distinct peak-like structure existing in Fig. 4(c) and (d).
@@
Fig. 5 presents the relaxation time of different particle species as functions of temperature at $\mu_{B}=0$ and $\mu_{B}=800$ MeV. For the $u,d$ isospin symmetric quark matter, the relaxation time $\tau_{u}=\tau_{d}$ and $\tau_{\bar{u}}=\tau_{\bar{d}}$ always hold. In particular, $\tau_{u}=\tau_{d}=\tau_{\bar{u}}=\tau_{\bar{d}}$ at vanishing chemical potential. In the case $\mu_{B}=0$, it can be seen that the relaxation time $\tau_{s}$ of strange quark is larger than that of $u$ ($d$) quark near the region of chiral and mott phase transitions. The chiral crossover transition temperature of $s$ quark is about $250$ MeV at $\mu_{B}=0$. We can see that $\tau_{s}$ gradually approaches $\tau_{u}$ at high temperature after the chiral restoration of $s$ quark, which is just the expected behavior under the SU(3) symmetry. At finite chemical potential, this figure shows that $\tau_{u}\neq\tau_{\bar{u}}$ and $\tau_{s}\neq\tau_{\bar{s}}$ due to the different distribution function of quark and antiquark. Numerically, the relaxation time of quarks are larger than those of antiquarks.
@@
Figure 5: Relaxation time of quarks and antiquarks as functions of temperature for $\mu_{B}=0$ and $\mu_{B}=800\,$MeV. The solid (hollow) triangles represent the $\eta/s$ at the point where these $\mu_{B}(T)$ line intersect the chiral crossover phase transition line (pion Mott transition line).
Fig. 5 also indicates that the relaxation for each particle species has a concave shape, which is to certain degree connected with the shear viscosity of QGP medium. The relaxation time describes the time scale required for a system to return to equilibrium after being subjected to a small perturbation. A shorter relaxation time indicates stronger interactions and more frequent collisions between particles. The shear viscosity coefficient is a macroscopic hydrodynamic quantity that describes a fluid’s ability to resist shear deformation, i.e., the relative motion between different fluid layers. A smaller viscosity coefficient means the fluid flows more “smoothly”. From the perspective of kinetic theory, the shear viscosity coefficient is proportional to the relaxation time. A short relaxation time implies frequent collisions between particles, leading to localized momentum that cannot be effectively transferred between fluid layers. Consequently, the fluid is more prone to shear deformation, meaning the viscosity coefficient is small. Therefore, in the phase diagram where the relaxation time has a concave shape, the viscosity coefficient probably also exhibit a minimum. In the next subsection we will see that the numerical results indeed exhibit this trend.
III.3 Shear viscosity under different conditions
@@
In the following, we elaborate on the shear viscosity of QGP under different conditions. First of all, we present in Fig. 6 the shear viscosity to entropy density ratio as functions of temperature at different chemical potentials. The dot, solid triangles and hollow triangles represent respectively the $\eta/s$ at the points where these $T(\mu_{B})$ lines intersect the CEP, chiral crossover phase transition line and pion Mott transition line. For $\mu_{B}$=0, the value of $\eta/s$ are very small when the temperature is close to and above the meson Mott phase transition. The minimum of $\eta/s$ is about 0.1, close to the KSS limit, which indicates that the QGP medium is a nearly “perfect” fluid in the very high energy collisions. The value of $\eta/s$ increases with the rising chemical potential. It can be observed from Fig. 6 that $\eta/s$ is larger than 1 in the critical region, and the value will continually increase at lower temperature and larger chemical potential.
@@
Figure 6: Shear visicosity to entropy density ratio $\eta/s$ as functions of temperature at different $\mu_{B}$ in the PNJL model. The dot, solid triangles and hollow triangles represent the $\eta/s$ at the points where these $\mu_{B}(T)$ lines intersect the CEP, chiral crossover phase transition line and pion Mott transition line, respectively.
@@
The minimum of $\eta/s$ at $\mu_{B}=0$ is closer to the Mott phase transition (meson dissociation) than the chiral phase transition of $u,d$ quark, and $\eta/s$ increases rapidly with the occurrence of hadronization. In Fig. 6 some values of $\eta/s$ below the chiral phase line are also preserved to display the variation of $\eta/s$ near the chiral crossover phase transition region. This figure clearly indicates that the properties of $\eta/s$ much depends on temperature, in particular near the phase transition. However we should keep in mind that the collisions between hadrons dominate the transport process in the hadronic phase at low temperature, which is not considered in the present work. The QGP phase in this study means the deconfined QCD medium within the PNJL model. It can be observed from Fig. 6 that the location of the minimum $\eta/s$ at finite chemical potential become more and more flat, and the relationship with phase transitions is not as distinct as that of $\mu_{B}=0$.
It is interesting to compare the PNJL results against those in the NJL model with a lower $T_{C}$. We present in Fig. 7 the temperature-dependent $\eta/s$ for a series of baryon chemical potential $\mu_{B}=0$, $400$, $800$, $948$ and $1200$ MeV in the NJL model. Compare Fig. 6 and Fig. 7, we can see that the two models give similar behavior of $\eta/s$. For small chemical potentials, the $\eta/s$ in the NJL model also exhibit a characteristic non-monotonic behavior with rising temperature, and the minimum appears at a temperature higher than the chiral crossover phase transition temperature (marked with the solid triangle). For the case of vanishing baryon chemical potential the minimum value of $\eta/s$ in the NJL model is slightly larger than that in the PNJL model, but are still close to 0.1. At the CEP, the values of $\eta/s$ are above 1 in both models, which indicates to some extent of the universality of shear viscosity near the critical region of strongly interacting matter.
Figure 7: Shear visicosity to entropy density ratio $\eta/s$ as functions of temperature at different $\mu_{B}$ in NJL model. The dot, solid triangles and hollow triangles represent the $\eta/s$ at the points where these $T(\mu_{B})$ lines intersect the CEP, chiral crossover phase transition line and pion Mott transition line, respectively.
To understand the behavior of $\eta/s$ in the $T-\mu_{B}$ plane in Fig. 6 derived in the PNJL model, we decompose the $\eta/s$ into shear viscosity coefficient $\eta$ and entropy density $s$. The scaled shear viscosity $\eta/T^{3}$ and entropy density $s/T^{3}$ are plotted in Figures 8 and 9. We can see that the $\eta/T^{3}$ for a given small chemical potential has a minimum with the increasing temperature and the corresponding $s/T^{3}$ only changes slightly at temperature higher than the Mott phase transition. Therefore, the existence of a minimum of $\eta/s$ is mainly attributed to the temperature-dependent behavior of $\eta/T^{3}$. Besides, below the chiral phase temperature, the value of $\eta/T^{3}$ increases quickly with the decrease of temperature, and meanwhile the value of $s/T^{3}$ declines sharply. They together lead to the quick enhancement of $\eta/s$ below the chiral phase transition temperature, as shown in Fig. 6.
Figure 8: The scaled shear viscosity $\eta/T^{3}$ as functions of temperature at different chemical potential $\mu_{B}=0$, $400$, $800$, $873$ and 12001200\,MeV in the PNJL model. The dot, solid triangles and hollow triangles represent the $\eta/T^{3}$ at the points where these $\mu_{B}(T)$ lines intersect the CEP, chiral crossover phase transition line and pion Mott transition line, respectively.
Figure 9: The scaled entropy density $s/T^{3}$ as function of temperature at different $\mu_{B}$ in the PNJL model.The dot, solid triangles and hollow triangles represent the $s/T^{3}$ at the points where these $\mu_{B}(T)$ lines intersect the CEP, chiral crossover phase transition line and pion Mott transition line, respectively.
@@
Fig. 10 describes the behaviors of $\eta/s$ as functions of chemical potential at several temperatures lower than or equal to 250250\,MeV. For the temperature higher than that of the chiral phase transition of $u(d)$ quark, e.g., $T=220$ or 250250\,MeV, it can be seen that $\eta/s$ monotonically increases with the rising chemical potential. The nonmonotonicity appears with the decrease of temperature, and the location of the minimum of $\eta/s$ gradually approaches to the chiral crossover phase transition (solid triangle in Fig. 10), particularly in the critical region.
Figure 10: $\eta/s$ as functions of baryon chemical potential at different temperatures in the PNJL model. The solid triangles, dots, and hollow triangles represent the $\eta/s$ at the points where these $T(\mu_{B})$ lines intersect the chiral corssover transition line, the first-order chiral phase transition line and pion Mott transition line, respectively. For T=50,75,100T=50,75,100\,MeV, the two dots on each isotherm correspond to respectively the high-density chiral restoration phase and low-density chiral breaking phase on the boundaries (binodal line) of the first-order phase transition.
@@
Fig. 10 also presents the values of $\eta/s$ at $T=100,75,50$ MeV where the first-order phase transition associated with a spinodal structure is involved. The solid dots on each isotherm with the same color indicates the value of $\eta/s$ on the boundaries of the first-order phase transition. From higher chemical potentials to lower ones, the first solid dot encountered on each curve corresponds to the high-density phase of the first-order chiral phase transition (in chiral restoration phase), and the second solid dot corresponds to the low-density phase of the chiral phase transition (in chiral breaking phase). This figure indicates that the values of $\eta/s$ are quite different in the chiral broken and chiral restored phases at lower temperature. The interval between the two dots on each isotherm corresponds to the metastable and unstable region.
Different from the first-order transition, the hollow triangles in Fig. 10 show that each $T(\mu_{B})$ line for T=100,75,50T=100,75,50\, MeV crosses only once the pion Mott phase transition line in the metastable or unstable phase. This can be understood in conjunction with Fig. 3. As illustrated in Fig. 3, the pion Mott transition line intersects solely with the high-density boundary of the first-order transition and then enters the metastable and unstable phases as temperature decreases. Since each point on the Mott line is defined by a specific $(T,\rho_{B})$ value which corresponds to a unique set of $(T,\mu_{B})$. Therefore the $\eta/s$ curve as a function of baryon chemical potential at a fixed temperature (e.g., T=100T=100\,MeV) in Fig. 10 intersects the Mott transition line at only one point.
Figure 11: $\eta/s$ as functions of net baryon density at different temperatures in the PNJL model. The solid triangles, dots, and hollow triangles represent the $\eta/s$ at the points where these $T(\rho_{B})$ lines intersect the chiral corssover transition line, the first-order chiral phase transition line and pion Mott transition line, respectively. For T=50,75,100T=50,75,100\,MeV, the two dots on each isotherm correspond to respectively the high-density chiral restoration phase and low-density chiral breaking phase on the binodal line of the first-order phase transition.
@@
Corresponding to Fig. 10, Fig. 11 presents the value of $\eta/s$ in the $T-\rho_{B}$ plane for different temperatures. The dashed curve connecting different isotherms is the boundary of the first-order phase transition. The behaviors of $\eta/s$ at $T=200$ MeV or higher temperatures are similar to those in the $T-\mu_{B}$ plane plotted in Fig. 10. However, compared to Fig. 10, in the $T-\rho_{B}$ plane the distance between the net baryon density of the minimum $\eta/s$ and that of the chiral crossover phase transition (marked with the solid triangle) appears to be enlarged when approaching the critical region of the CEP from the high-temperature side. The reason is that the net baryon density is sensitive to chemical potential in the critical region. Small variation of chemical potential will lead to a big change of net baryon density. For the first-order phase transition, the difference of $\eta/s$ in the two phases is small near the CEP, since the dynaical quark masses of quarks are slight different. However, it can be seen that at lower temperature, e.g., $T=50$ MeV, $\eta/s$ in the chiral restored QGP phase is much larger than the chiral breaking phase. It also indicates the shear viscosity to entropy density ratio near the first-order phase transition at lower temperatures very depends on the net baryon density.
@@
Figure 12: The value of $\eta/s$ on the phase boundaries including the chiral corssover and first-order phase transitions in the PNJL model.
Fig. 10 and Fig. 11 also present that, in high-density region, the value of $\eta/s$ increases with the decline of temperature, which is qualitatively consistent with the variation of shear viscosity of common liquids. Importantly, the change of $\eta/s$ with the physical conditions inevitably causes us to pay more attention to the viscosity of QGP medium at finite temperature and density. The value of $\eta/s$ is crucial in diagnosing the QCD phase transition signals in the hydrodynamical simulation of heavy-ion collisions with colliding energies at several GeV, close to the lower side of the beam energy scan at RHIC STAR. that $\eta/s\approx 0.65\pm 0.15$ in Au+Au reactions at $\sqrt{s}=2.4$ GeV [76].